OpenStudy (anonymous):

1/4(x+5)- 5/3= 1/3(x-2)

OpenStudy (anonymous):

\[\frac{ 1 }{ 4(x+5) }-\frac{ 5 }{ 3 }=\frac{ 1 }{ 3(x-2) }\]

OpenStudy (anonymous):

u there?